Solve the system. (Give your answer as an improper fraction. If the system has infinitely many solutions, express your answer in terms of k, where t = t(k) and v = k. If the system has no solution, enter NONE for each answer.)
1/2t-1/5v=5/2
2/3t+1/4v=29/12

Answer by stanbon(75887)

1/2t-1/5v=5/2
2/3t+1/4v=29/12
-----
Multiply thru 1st by 10
Multiply thru 2nd by 12
-----
5t - 2v = 25
8t + 3v = 29
------
Multiply thru 1st by 3
Multiply thru 2nd by 2
-----
15t - 6t = 3*25
16t + 6t = 2*29
----
Add and solve for "t":
31t = 133
t = 133/13
-----
Substitute into 5t-2v = 25 to solve for "v":
5(133/13)-2v = 25
2v = 340/13
v = 170/13
=======================
Ans: The system is independent.
=======================
Cheers,
Stan H.